OpenStudy (anonymous):

Jorge plans to paint a bedrrom set wallthat is shaped like a trapezoid.the bottom edge of the wall is 22.5 feet long and the top edge of the wall is 9.5 feet long if the wall is 8 feet tall what is the area of the wall..>:)

jimthompson5910 (jim_thompson5910):



OpenStudy (anonymous):

okk

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

base1 = 9.5 b1 = 9.5

jimthompson5910 (jim_thompson5910):

base2 = 22.5 b2 = 22.5

jimthompson5910 (jim_thompson5910):

height = 8 h = 8

jimthompson5910 (jim_thompson5910):

so we'll be using b1 = 9.5 b2 = 22.5 h = 8

jimthompson5910 (jim_thompson5910):

in the formula A = h*(b1+b2)/2

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

so what do you get when you do this?

jimthompson5910 (jim_thompson5910):

A = h*(b1+b2)/2 A = 8*(9.5 + 22.5)/2 A = ???

OpenStudy (anonymous):

im cufused here

jimthompson5910 (jim_thompson5910):

where are you stuck

OpenStudy (anonymous):

i don't relly know

OpenStudy (anonymous):

i jsut cufused ont h hwolet thing here

jimthompson5910 (jim_thompson5910):

at this point, you just use your calculator to evaluate 8*(9.5 + 22.5)/2

jimthompson5910 (jim_thompson5910):

do you have a calculator?

OpenStudy (anonymous):

yes one sec k i will di that part now and tlel what i get.>:) K

jimthompson5910 (jim_thompson5910):

alright